How the volume occupied by 1 mole atom of diatomic gas at NTP will be 11.2 litres?

Grade:11

6 Answers

ABHISHEK JAIN AskiitiansExpert-IITD
13 years ago

Dear Chanchal,

  Mole = Volume/22.4 (At S.T.P or N.T.P.)

so according to this formula 1 Mole of gas occupies 22.4 litres at N.T.P.

 

suuppose diatomic gas is O2

 1 mole of O2 gas occupies  22.4 lit at N.T.P.

since O2 consist of two atoms means 1 mole of O2 contains 2 mole atom of oxygen.

thus we can say since 1 mole of O2 or 2 mole atoms of oxygen occupies = 22.4 litres at N.T.P.

therefore 1 mole atom of oxygen occupies = 22.4/2 = 11.2 litres at N.T.P.

 

We are all IITians and here to help you in your IIT JEE preparation.

 

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

YOUR NAME

ABHISHEK JAIN

AskiitiansExpert-IITD

 

Rohit Singh
13 years ago

Its like that.

1 mole of diatomic gas at STP or NTP=22.4L

this means 2*(i mole of atom not molecule)=22.4

=> 1 mole atom(not molecule)=22.4/2

=11.2 litre.

suresh
9 years ago
why one mole gas is having same volume at stp and ntp, both are not same temperatures. according to charles law, v is proportionl to volume.
suresh
9 years ago
why one mole gas is having same volume at stp and ntp, both are not same temperatures. according to charles law, v is proportionl to temperature
Bhawani shankar Sahu
7 years ago
actually the volume at NTP and STP are not equal. at STP the volume of 1 mole of gas is 22.4 lt but at NTP the volume is 24.02 lt.  you can derive this from PV=nRT
Yash Chourasiya
3 years ago
Hello Student

We know that, 1 mole of gas occupies 22.4 L at NTP.
1 mole of diatomic gas (O2) gas occupies 22.4 L at NTP
O2 contains 2 moles of oxygen atom.
Hence 2 moles of oxygen atom occupies 22.4 L at NTP
Mole = Volume / 22.4
Therefore, 1 mole of oxygen atom contains 22.4 / 2 = 11.2 L at NTP.
Which means the volume occupied by 1 mole atom of diatomic gas at NTP will be 11.2 litres.

I hope this answer will help you.